OpenStudy (mary1999):

enter an ordered pair corresponding to a 90 degree counterclockwise rotation of the point (14,10) about the origin followed by the reflection the line x=6

OpenStudy (mathmate):

Hint: \(R_{90}(x,y)=(-y,x)\) For reflection about x=k, you would translate to the reflection plane, reflect, then translate back. so we need the composite transformation: \(S_{x=k}(x,y)=(x+k,y)o(-x,y)o(x-k,y)\) which reduces to \(S_{x=k}\(x,y)=(-(x-k)+k,y)=(2k-x,y))

OpenStudy (mary1999):

i'm sorry but that really confused me